Find the concentration of holes and electrons in a p-type germanium at 300k if the conductivity is 100ohm-cm, atoms 10^3=44×10^3 n1¹= 1800cm²/vc


1
Expert's answer
2021-05-13T10:28:25-0400

"T=300K, \sigma= 100\Omega - cm,\mu_p=1800cm^2/vc"


As we know,


"\sigma=n_1 e\mu_p"


Concentration of holes-

"n_1 \text{ or } p=\dfrac{\sigma}{e\mu_o}"


    "=\dfrac{100}{1.6\times 10^{-19}\times 1800}=3.468\times 10^{17} cm^3"



Concentration of electron


"e^{-1}=n=\dfrac{n_i^2}{p}\\[9pt]=\dfrac{(25.\times 10^{13})^2}{3.46\times 10^{17}}\\[9pt]=1.802\times 10^{9}/ cm^3"